Noy is twice as old as Kris. In eight years, Noy will be 4 years older than Kris. Find their ages.

Answer by Cromlix(4381)

Noy is twice as old as Kris
N = 2K
In eight years Noy will be
4 years older than Kris.
N + 8 = (K + 8) + 4
Replace N with 2K from 1st equation.
2K + 8 = K + 12
Collect like terms
2K - K = 12 - 8
K = 4
Current ages:
Kris = 4 years old
Noy = 8 years old.
Hope this helps.
:-)